The method of tree ring dating gave the following years A.D. for an archaeological excavation site. Assume that the population of x values has an approximately normal distribution.

1292 1264 1229 1243 1268 1316 1275 1317 1275

(a) Use a calculator with mean and standard deviation keys to find the sample mean year x and sample standard deviation s. (Round your answers to the nearest whole number.)

x = A.D.
s = yr


(b) Find a 90% confidence interval for the mean of all tree ring dates from this archaeological site. (Round your answers to the nearest whole number.)

lower limit     A.D.
upper limit     A.D.
Answer #1

a)

x =1275
s = 30

b)


sample mean, xbar = 1275
sample standard deviation, s = 30
sample size, n = 9
degrees of freedom, df = n - 1 = 8

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, tc = t(α/2, df) = 1.86



CI = (1275 - 1.86 * 30/sqrt(9) , 1275 + 1.86 * 30/sqrt(9))
CI = (1256 , 1294)


lower limit = 1256
Upper limit = 1294